# Laser mode-hopping assisted all-optical single beam pulsed atomic magnetometer

### v1

@@ V Conclusion
We have demonstrated a single-beam all-optical pulsed atomic magnetometer enhanced by laser mode-hopping. Optical pumping assisted by laser mode-hopping, which leads to a discontinuous pumping profile, effectively enhances the transverse spin polarization. Furthermore, mode-hopping offers a convenient method for setting probing stage beam off-resonant which suppress the absorption and maintain large FID signal. We have reached a sensitivity of 3.77 pT/Hz1/2 which is sufficent for applications of magnetic anomaly detection in earth’s magnetic field. The sensitivity of our magnetometer seems to be limited by the performance of the frequency counter. In order to estimate the potentially achievable sensitivity, CRLB due to the non-magnetic oriented noise was calculated, which was 191 $\textrm{fT}/\sqrt{\textrm{Hz}}$ in our experimental condition. Considering the calculated CRLB, we expect that our magnetometer can reach a sensitivity of a sub pT/Hz1/2 which is similar to the sensitivity limited by photon shot-noise. A laser cavity optimization considering laser mode-hopping in fabrication of a DBR LD may further improve the sensitivity by ehancing optical pumping. Our approach based on laser mode-hopping can be applied to miniaturization of all-optical atomic magnetometers with sub-pT/Hz1/2 sensitivities.
@@ VII Appendix
VII.1 Consideration on Faraday Rotation of Elliptically Polarized Beam
The spin polarization formed assisted by mode-hopping enters the probing stage where elliptically polarized light undergoes optical Faraday rotation, then measured by a polarimetry. In a linear polarization basis ($\hat{e}_{H},\hat{e}_{V}$), the elliptical polarization state is represented as $(\cos(\theta),\sin(\theta)\exp(i\phi))$. Here, $\theta$ is the tilt angle, and $\phi$ is the retardation angle. The change in polarization state of the beam after passing through the atomic vapor can be expressed by the matrix $M_{c}=\begin{pmatrix}{1+e^{i\Phi}}&{i(1-e^{i\Phi})}\\ {i(e^{i\Phi}-1)}&{1+e^{i\Phi}}\end{pmatrix}$
where $\Phi=|s|\Phi_{0}(t)$ is the Faraday rotation angle. Here, $\Phi_{0}(t)$ is damped oscillation in our experiment and $|s|$ is the size of photon spin vector which is calculated by $\sin(2\theta)\sin(\phi)$. The matrix representation of the HWP is $M_{h}=\begin{pmatrix}{\cos(2\theta_{h})}&{\sin(2\theta_{h})}\\ {\sin(2\theta_{h})}&{-\cos(2\theta_{h})}\end{pmatrix}$, so the final polarization state of the elliptically polarized beam measured by the polarimeter is $M_{h}\cdot M_{c}\cdot(\cos(\theta),\sin(\theta)\exp(i\phi))$. The balanced signal is then calculated by the difference between squares of each orthogonal linear polarization components of the beam: $4I_{0}[\cos(4\theta_{h}+\Phi)\cos(2\theta)+\cos(\phi)\sin(4\theta_{h}+\Phi)\sin(2\theta)]$. Here, $I_{0}$ is the intensity of the beam after absorption throughout the vapor cell. When $1\gg\Phi$, the oscillating signal simplifies to
@@ VII.1 Consideration on Faraday Rotation of Elliptically Polarized Beam
Figure 7: The polarization state of incident beam(upper plot) and relative FID amplitude(bottom plot) as a function of QWP angle. When the QWP angle is 25 where relative FID amplitude is maximized, $\theta$ is 30 and $\phi$ is 129 . The red line shows $\sin(4\theta_{q})$ behavior to the QWP angle $\theta_{q}$. Blue line illustrates Eq. (8) calculated based on the polarization state on the top plot.
By adjusting the HWP angle $\theta_{h}$, the baseline component of the balanced signal can be eliminated. The amplitude coefficient proportional to $\Phi_{0}(t)$, simultaneously, is maximized.
Previous study has shown that signal amplitude in single-beam SERF magnetometers follows a $\sin(4\theta_{q})$-dependence on QWP angle $\theta_{q}$shah2009spin . Our all-optical single beam pulsed magnetometer shows similar behavior under the bias field of 14.3 $\mu T$ which is comparable to the Earth’s ambient magnetic field. The top plot of Figure. 7 shows the polarization state of the beam before entering the vapor cell as measured by a polarimeter(Thorlabs PAX 5710IR1) at various QWP angles. The bottom plot shows the relative FID amplitude measured at different QWP angles. The red line illustrates $|\sin(4\theta_{q})|$ behavior, while the blue line represents equation (8), which is calculated based on the polarization state shown in the upper plot. HWP angle $\theta_{h}$ is adjusted to compensate signal baseline to be zero for each measurement.
We determined the optimal QWP angle for single-beam all-optical magnetometer to be 25 with a 133 retardation, suggesting an elliptically polarized beam as the most effective polarization state. The elliptical polarization strikes a balance between the most efficient pumping of circular polarization and the most effective Faraday rotation that occurs with linear polarization. Note that for FID measurements, compensating baseline through HWP angle adjustment is essential. Without baseline compensation, the FID amplitude may decrease, and if the beam intensity is too high, the baseline component may saturate the entire signal. To avoid signal saturation on the BPD signal, in our measurement, the beam intensity in front of the vapor cell is kept at 1.43 mW during this measurement.

### v2

@@ V Conclusion
We have demonstrated a single-beam all-optical pulsed atomic magnetometer enhanced by laser mode-hopping. When optical pumping is assisted by laser mode-hopping, it leads to a discontinuous pumping profile, effectively enhancing the transverse spin polarization. Furthermore, mode-hopping offers a convenient method for setting the probing stage beam off-resonant, which suppresses absorption and maintains a large FID signal. We achieved a sensitivity of 0.6 pT/Hz1/2 which is sufficient for applications to detect magnetic anomalies in earth’s magnetic field. The sensitivity of our magnetometer seems to be limited by photon-shot-noise and the magnetic field noise induced by the current noise in the current supply for driving the bias magnetic field coil. In order to estimate the potentially achievable sensitivity, the CRLB due to the non-magnetic oriented noise was calculated, which was 131 $\textrm{fT}/\sqrt{\textrm{Hz}}$ in our experimental condition. Considering the calculated CRLB, we expect that our magnetometer can reach a sensitivity of a sub pT/Hz1/2 which is similar to sensitivity limited by photon shot-noise. Laser cavity optimization considering laser mode-hopping during the fabrication of a DBR LD may further improve sensitivity by enhancing optical pumping. Our approach based on laser mode-hopping can be applied for the miniaturization of all-optical atomic magnetometers with sub-pT/Hz1/2 sensitivities.
@@ VII Appendix
VII.1 Consideration of Faraday Rotation of Elliptically Polarized Beam
The spin polarization formed assisted by mode-hopping enters the probing stage where elliptically polarized light undergoes optical Faraday rotation, then measured by a polarimetry. With a linear polarization basis ($\hat{e}_{H},\hat{e}_{V}$), the elliptical polarization state is represented as $(\cos(\theta),\sin(\theta)\exp(i\phi))$. Here, $\theta$ is the tilt angle, and $\phi$ is the retardation angle. The change in polarization state of the beam after passing through the atomic vapor can be expressed by the matrix $M_{c}=\begin{pmatrix}{1+e^{i\Phi}}&{i(1-e^{i\Phi})}\\ {i(e^{i\Phi}-1)}&{1+e^{i\Phi}}\end{pmatrix}$
where $\Phi=|s|\Phi_{0}(t)$ is the Faraday rotation angle. Here, $\Phi_{0}(t)$ is damped oscillation in our experiment and $|s|$ is the size of the photon spin vector which is calculated by $\sin(2\theta)\sin(\phi)$. The matrix representation of the HWP is $M_{h}=\begin{pmatrix}{\cos(2\theta_{h})}&{\sin(2\theta_{h})}\\ {\sin(2\theta_{h})}&{-\cos(2\theta_{h})}\end{pmatrix}$, so the final polarization state of the elliptically polarized beam measured by the polarimeter is $M_{h}\cdot M_{c}\cdot(\cos(\theta),\sin(\theta)\exp(i\phi))$. The balanced signal is then calculated by the difference between the squares of each orthogonal linear polarization components of the beam: $4I_{0}[\cos(4\theta_{h}+\Phi)\cos(2\theta)+\cos(\phi)\sin(4\theta_{h}+\Phi)\sin(2\theta)]$. Here, $I_{0}$ is the intensity of the beam after absorption throughout the vapor cell. When $1\gg\Phi$, the oscillating signal simplifies to
@@ VII.1 Consideration of Faraday Rotation of Elliptically Polarized Beam
Figure 7: The polarization state of the incident beam (upper plot) and relative FID amplitude (bottom plot) as a function of QWP angle. When the QWP angle is 25 where the relative FID amplitude is maximized, $\theta$ is 30 and $\phi$ is 129 . The red line shows $\sin(4\theta_{q})$ behavior to the QWP angle $\theta_{q}$. The blue line illustrates Eq. (8) calculated based on the polarization state on the top plot.
By adjusting the HWP angle $\theta_{h}$, the baseline component of the balanced signal can be eliminated. The amplitude is coefficient proportional to $\Phi_{0}(t)$, and is maximized at the same time.
A previous study has shown that signal amplitude in single-beam SERF magnetometers follows a $\sin(4\theta_{q})$-dependence on the QWP angle $\theta_{q}$shah2009spin . Our all-optical single beam pulsed magnetometer shows similar behavior under the bias field of 14.3 $\mu T$ which is comparable to the Earth’s ambient magnetic field. The top plot of Figure. 7 shows the polarization state of the beam before entering the vapor cell as measured by a polarimeter (PAX 5710IR1, Thorlabs) at various QWP angles. The bottom plot shows the relative FID amplitude measured at different QWP angles. The red line illustrates $|\sin(4\theta_{q})|$ behavior, while the blue line represents equation (8), which is calculated based on the polarization state shown in the upper plot. The HWP angle $\theta_{h}$ is adjusted to compensate so the signal baseline will be zero for each measurement.
We determined the optimal QWP angle for the single-beam all-optical magnetometer to be 25 with a retardation of 133 , suggesting an elliptically polarized beam as the most effective polarization state. The elliptical polarization strikes a balance between the most efficient pumping of circular polarization, and the most effective Faraday rotation that occurs with linear polarization. Note that baseline compensation by adjusting the HWP angle is essential for FID measurements. Without baseline compensation, the FID amplitude may decrease, and if the beam intensity is too high, the baseline component may saturate the entire signal. To avoid signal saturation on the BPD signal, in our measurement, the beam intensity in front of the vapor cell is kept at 1.43 mW during this measurement.